# Lifetime-Limited Linewidth Measurements of the 3C and 3D Soft X-ray Transitions in Ni XIX

### v1

@@ III Results and Discussions
Line 3C | Line 3D | |||
Experiment | Theory | Experiment | Theory | |
20.50 1.70 | 20.867 0.026 | 7.70 0.70 | 7.570 0.027 | |
2.17 0.18 | 2.204 0.003 | 0.84 0.08 | 0.828 0.003 | |
3.12 0.26 | 3.170 0.004 | 1.17 0.11 | 1.150 0.004 | |
32.11 2.68 | 31.54 0.040 | 85.48 7.84 | 86.95 0.310 |
@@
Figure 2 compares our results with earlier experimental data, observations, and predictions. Our own calculations using large-scale configuration interaction (large-CI) Cheung et al. (2021) and FAC-CI Gu (2008) methods, both accounting for the relativistic Breit interaction and quantum electrodynamics (QED) effects, yield an oscillator-strength ratio which agrees with our experiment within the uncertainties. Details of large-CI computations are given in Appendix A. Additional comparisons with other theoretical values from the literature showed typical departures within a 1–2 range, with the largest one from Zhang and Sampson (1989).
@@
Three laboratory measurements of the 3C/3D intensity ratio of Ni XIX based on electron-impact excitation (EIE) were reported Brown et al. (2001c); Gu et al. (2004); Chen et al. (2006). The collisional excitation ratios from all three previous experiments span 1.90–2.35, slightly lower than the oscillator strength ratio measured in our experiment, but all four experiments are mutually consistent within uncertainties. Observations of the solar corona Hutcheon et al. (1976); McKenzie et al. (1980) and Capella Behar et al. (2001) show some scatter in the 3C/3D ratio, but in the absence of uncertainty estimates we cannot quantify the level of agreement with laboratory measurements.
@@
Figure 3 (a) shows a comparison between measured 3C and 3D natural linewidths obtained using two different methods, which agree (see dashed diagonal lines) with each other and our predictions. Both our calculations using large-CI and FAC-CI methods agree within the 1 uncertainty of our experimental results. As can be seen in Figure 3 (b), the other older calculations from Zhang and Sampson (1989) and Safronova et al. (2001) show disagreement, however, their values just fall outside of the 2-sigma ellipse of mean natural linewidths, whereas somewhat recent calculations presented in Refs. Dong et al. (2003); Jönsson et al. (2014); Santana et al. (2015) show agreement within the 1-sigma ellipse.
@@
The individual oscillator strengths of 3C and 3D can be derived from the natural linewidth using the relation , where . and represent the statistical weights of the initial () and final states (), respectively. The is the transition wavelength given in nanometers taken from Gu et al. (2007), and is the reduced Planck constant. Moreover, the experimental lifetime and transition rates can also be derived from the natural linewidth measurements. All of these measured quantities agree very well with our predictions, given in Tab. 2.
@@
Our present experimental validation of oscillator strengths effectively eliminates uncertainties in atomic data as potential contributors to the observed discrepancies in nickel opacity measurements Nagayama et al. (2019) and iron opacity Bailey et al. (2007); Fontes et al. (2015); Kühn et al. (2022). The measured oscillator strengths offer direct application in astrophysical modeling, enabling the diagnosis of turbulent velocity in moderately optically thick plasmas Xu et al. (2002b); Werner et al. (2009). Upon comparing with oscillator strength ratios and natural linewidths from established databases such as NIST-ASD Kramida et al. (2022), AtomDB Foster et al. (2012), Chianti Del Zanna and Ishikawa (2009), and SPEX Kaastra et al. (1996), significant discrepancies were identified with our experimental results. There is a pressing need for updating these databases with such precise measurements in view of the crucial role of accurate data in modeling observational spectra from existing missions like Chandra and XMM-Newton, as well as upcoming observations with XRISM Tashiro et al. (2018), which has recently been successfully launched, as well as future planned missions such as Athena (Pajot et al., 2018), LEM (Kraft et al., 2022), HUBS (Cui et al., 2020), Arcus (Heilmann et al., 2022), HiReX (Nicastro et al., 2021), and Lynx (Schwartz et al., 2019). The present combination of experimental benchmark and converged calculations is a crucial consistency check for atomic data required to interpret X-ray observations in astrophysics, fusion, and high-energy density plasma research.
@@ Acknowledgements.
This research was funded by the Max Planck Society (MPG) and the German Federal Ministry of Education and Research (BMBF) under project 05K13SJ2. C.S. acknowledges support from MPG and NASA under grant number 80GSFC21M0002. F.S.P. and M.A.L. acknowledge support from the NASA Astrophysics Program. The theoretical work has been supported by the US NSF Grant No. PHY-2012068, No. PHY-2309254, US Office of Naval Research Grant No. N00014-20-1-2513, and by the European Research Council (ERC) under the Horizon 2020 Research and Innovation Programme of the European Union (Grant Agreement No. 856415). Calculations were performed in part using the computing resources at the University of Delaware, in particular the Caviness and DARWIN high-performance computing clusters. We acknowledge DESY (Hamburg, Germany), a member of the Helmholtz Association HGF, for the provision of experimental facilities. Parts of this research were carried out at PETRA III. We thank Jens Viefhaus and Rolf Follath for valuable discussions on X-ray monochromator resolution and performance, and the synchrotron-operation team and P04 team at PETRA III for their skillful and reliable work. MSS thanks MPIK, Heidelberg for hospitality.*
@@ Appendix A Theory
The calculations are carried out using a large-scale configuration interaction (CI) method Cheung et al. (2021), including correlations from all ten electrons. Basis sets of increasing sizes are used to check for convergence of the values. The basis set is designated by the highest principal quantum number for each partial wave included. For example, [] means that all orbitals up to are included for the partial waves. We begin by considering all possible single and double excitations to orbitals up to from the , even and , , odd configurations, correlating 8 electrons. We verified that inclusion of the even and , odd configurations as basic configurations has a negligible effect on the energies and relevant matrix elements. The calculated contributions to the energies of Ni XIX are listed in Tab. 3. The results are compared with a revised analysis of the experimental data Kramida (2019). We use -coupling and NIST-style LS-coupling term designations for comparisons. Contributions to the reduced matrix elements and and the oscillator strengths ratios are listed in Table 4. The energy ratio is 1.018 and the oscillator strength ratio is 2.64(2).
@@
To assess the impact of triple excitations, we consider a broad range of configurations up to . As demonstrated in Tables 3 and 4, these excitations result in negligible corrections to both energies and matrix elements. Subsequently, we expand the basis set to , leading to a significant improvement in the agreement of energies with experimental values and a minimal shift in the ratio (-0.008). Further accounting of contributions from the shell with the basis improves agreement with experimental energies, albeit with a marginal contribution to the ratio (-0.006). A comparison of results for and obtained in length and velocity gauges reveals only a marginal difference of 0.001 for the basis. Expanding the basis set to produces a modest improvement in energies compared to experiment, accompanied by a slight shift in the ratio (-0.001). Further expansion to results in minor corrections to energies, with an even smaller contribution to the ratio (-0.0005). The quantum electrodynamic (QED) contributions are included following Ref. Tupitsyn et al. (2016). The QED inclusion has small effect on the individual line energies, however, a considerable contribution to the energy difference of 3C and 3D, see Tab. 3. Furthermore, the ratio is changed by -0.01 by accounting for QED.
@@
Additionally, we compute the transition rates for all other transitions contributing to the radiative decay of the 3 and 3 levels. The sums of these rates are small and listed in Tab. 4. Linewidth values correspond to total transition rates, with uncertainties computed from the largest uncertainties in the 3C and 3D matrix elements, including additional configurations in CI space and QED. Additional uncertainties due to orbitals for transition rates and linewidths, based on recent studies of Fe16+ Shah et al. (2024), are included, and the final uncertainties are computed based on the relative difference.

### v2

@@ III Results and Discussions
Figure 2 compares our results with earlier experimental data, observations, and predictions. Our own calculations using large-scale configuration interaction (large-CI) Cheung et al. (2021) and FAC-CI Gu (2008) methods, both accounting for the relativistic Breit interaction and quantum electrodynamics (QED) effects, yield an oscillator-strength ratio which agrees with our experiment within the uncertainties. Details of large-CI computations are given in the Appendix A. Additional comparisons with other theoretical values from the literature showed typical departures within a 1–2 range, with the largest one being from Zhang and Sampson (1989).
@@
Three laboratory measurements of the 3C/3D intensity ratio of Ni XIX based on electron-impact excitation were reported Brown et al. (2001c); Gu et al. (2004); Chen et al. (2006). The collisional excitation ratios from all three previous experiments span 1.90–2.35, slightly lower than the oscillator strength ratio measured in our experiment, but all four experiments are mutually consistent within uncertainties. Observations of the solar corona Hutcheon et al. (1976); McKenzie et al. (1980) and Capella Behar et al. (2001) show some scatter in the 3C/3D ratio, but in the absence of uncertainty estimates we cannot quantify the level of agreement with laboratory measurements.
@@
Figure 3 (a) shows a comparison between measured 3C and 3D natural linewidths obtained using two different methods, which agree (see dashed diagonal lines) with each other and our predictions. Our calculations using both large-CI and FAC-CI methods agree within the 1 uncertainty of our experimental results. As can be seen in Figure 3 (b), the other older calculations from Zhang and Sampson (1989) and Safronova et al. (2001) show disagreement; however, their values fall just outside of the 2-sigma ellipse of mean natural linewidths, whereas somewhat recent calculations presented in Refs. Dong et al. (2003); Jönsson et al. (2014); Santana et al. (2015) show agreement within the 1-sigma ellipse.
@@
The individual oscillator strengths of 3C and 3D can be derived from the natural linewidth using the relation , where . and represent the statistical weights of the initial () and final () states, respectively. is the transition wavelength given in nanometers taken from Gu et al. (2007), and is the reduced Planck constant. The experimental oscillator strengths for 3C and 3D are found to be 2.17 0.18 and 0.84 0.08, respectively. This corresponds to lifetimes of 32.11 2.68 fs and 85.48 7.84 fs for 3C and 3D, respectively. All of these measured quantities agree very well with our predictions.
@@
Our present experimental validation of oscillator strengths effectively eliminates uncertainties in atomic data as potential contributors to the observed discrepancies in nickel opacity measurements Nagayama et al. (2019) and iron opacity Bailey et al. (2007); Fontes et al. (2015); Kühn et al. (2022). The measured oscillator strengths offer direct application in astrophysical modeling, enabling the diagnosis of turbulent velocity in moderately optically thick plasmas Xu et al. (2002b); Werner et al. (2009). Following a comparing with oscillator strength ratios and natural linewidths from established databases such as NIST-ASD Kramida et al. (2022), AtomDB Foster et al. (2012), Chianti Del Zanna and Ishikawa (2009), and SPEX Kaastra et al. (1996), significant discrepancies were identified with our experimental results. There is a pressing need to update these databases with such precise measurements in view of the crucial role of accurate data in modeling observational spectra from existing missions like Chandra and XMM-Newton, as well as upcoming observations with XRISM Tashiro et al. (2018), which was recently successfully launched, as well as future planned missions such as Athena (Pajot et al., 2018), LEM (Kraft et al., 2022), HUBS (Cui et al., 2020), Arcus (Heilmann et al., 2022), HiReX (Nicastro et al., 2021), and Lynx (Schwartz et al., 2019). The present combination of experimental benchmark and converged calculations is a crucial consistency check for atomic data required to interpret X-ray observations in astrophysics, fusion, and high-energy density plasma research.
@@ Acknowledgements.
This research was funded by the Max Planck Society (MPG) and the German Federal Ministry of Education and Research (BMBF) under Project No. 05K13SJ2. C.S. acknowledges support from MPG and NASA. F.S.P. and M.A.L. acknowledge support from the NASA Astrophysics Program. The theoretical work was supported by the US NSF Grant No. PHY-2012068 and No. PHY-2309254, U.S. Office of Naval Research Grant No. N00014-20-1-2513, and the European Research Council (ERC) under the Horizon 2020 Research and Innovation Programme of the European Union (Grant Agreement No. 856415). Calculations were performed in part using the computing resources at the University of Delaware, in particular the Caviness and DARWIN high-performance computing clusters. This theoretical work, including M.G.K.’s contributions, was carried out as part of a scientific collaboration during the calendar year 2020. We acknowledge DESY (Hamburg, Germany), a member of the Helmholtz Association HGF, for the provision of experimental facilities. Parts of this research were carried out at PETRA III. We thank Jens Viefhaus and Rolf Follath for valuable discussions on X-ray monochromator resolution and performance, and the synchrotron-operation team and P04 team at PETRA III for their skillful and reliable work. MSS thanks MPIK, Heidelberg, for hospitality.*
@@ Appendix A Theory
The calculations are carried out using a large-scale configuration interaction (CI) method Cheung et al. (2021), including correlations from all 10 electrons. Basis sets of increasing sizes are used to check for convergence of the values. The basis set is designated by the highest principal quantum number for each partial wave included. For example, [] means that all orbitals up to are included for the partial waves. We begin by considering all possible single and double excitations to orbitals up to from the and even configurations and , , and odd configurations, correlating 8 electrons. We verified that inclusion of the even and , odd configurations as basic configurations has a negligible effect on the energies and relevant matrix elements. The calculated contributions to the energies of Ni XIX are listed in Tab. 2. The results are compared with a revised analysis of the experimental data Kramida (2019). We use -coupling and NIST-style LS-coupling term designations for comparisons. Contributions to the reduced matrix elements and and the oscillator strengths ratios are listed in Table 3. The energy ratio is 1.018 and the oscillator strength ratio is 2.64(2).
@@
To assess the impact of triple excitations, we consider a broad range of configurations up to . As demonstrated in Tables 2 and 3, these excitations result in negligible corrections to both energies and matrix elements. Subsequently, we expand the basis set to , leading to a significant improvement in the agreement of energies with experimental values and a minimal shift in the ratio (). Further accounting of contributions from the shell with the basis improves agreement with experimental energies, albeit with a marginal contribution to the ratio (). A comparison of results for and obtained in length and velocity gauges reveals only a marginal difference of 0.001 for the basis. Expanding the basis set to produces a modest improvement in energies compared to experiment, accompanied by a slight shift in the ratio (). Further expansion to results in minor corrections to energies, with an even smaller contribution to the ratio (). The quantum electrodynamic (QED) contributions are included following Ref. Tupitsyn et al. (2016). The inclusion of QED has a small effect on the individual line energies, however, a considerable contribution to the energy difference of 3C and 3D, see Tab. 2. Furthermore, the ratio is changed by by accounting for QED.
@@
Additionally, we compute the transition rates for all other transitions contributing to the radiative decay of the 3 and 3 levels. The sums of these rates are small and are listed in Tab. 3. Linewidth values correspond to total transition rates, with uncertainties computed from the largest uncertainties in the 3C and 3D matrix elements, including additional configurations in CI space and QED. Additional uncertainties due to orbitals for transition rates and linewidths, based on recent studies of Fe16+ Shah et al. (2024), are included, and the final uncertainties are computed based on the relative difference.